OpenStudy (anonymous):

how do you put y=2/3x+7 in standard form?

OpenStudy (anonymous):

the std form is ax + by = c y = 2/3x -7 multiply both sides by 3 3y = 2x - 21 2x - 3y = 21

OpenStudy (anonymous):

but its positive 7

OpenStudy (anonymous):

Standard form is written like this: Ax + By + C = 0, where A and B are some coefficients of x and y, and C is a constant. Here we are given the slope intercept form, y = 2/3x + 7. We arrange it so everything is on a single side. -2/3x + y = 7 --> I brought the 2/3x over to the other side -2/3x + y - 7 = 0 --> I brought the same 7 over to the other side 2/3x - y + 7 = 0 --> I multiplied both sides of the equation by -1, 0*1 remains 0, and multiplying the left side of the equation by negative -1 switches the signs of the terms so it looks nice and clean. Although you could leave it at -2/3x + y - 7 = 0, I just multiplied both sides by -1 because I like my first coefficient to be positive. Therefore, the standard form is: 2/3x - y + 7 = 0 or -2/3x + y - 7 = 0

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

^They got it :) haha I messed up cause of the whole positive/ negative thing

OpenStudy (anonymous):

since when is there (=0)?

OpenStudy (anonymous):

i got 2x+3y=21